# ScreenMark: Watermarking Arbitrary Visual Content on Screen

### v2

@@ Abstract
Digital watermarking has demonstrated its effectiveness in protecting multimedia content. However, existing watermarking are predominantly tailored for specific media types, rendering them less effective for the protection of content displayed on computer screens, which is often multimodal and dynamic. Visual Screen Content (VSC), is particularly susceptible to theft and leakage via screenshots, a vulnerability that current watermarking methods fail to adequately address. To tackle these challenges, we propose ScreenMark, a robust and practical watermarking method designed specifically for arbitrary VSC protection. ScreenMark utilizes a three-stage progressive watermarking framework. Initially, inspired by diffusion principles, we initialize the mutual transformation between regular watermark information and irregular watermark patterns. Subsequently, these patterns are integrated with screen content using a pre-multiplication alpha blending technique, supported by a pre-trained screen decoder for accurate watermark retrieval. The progressively complex distorter enhances the robustness of the watermark in real-world screenshot scenarios. Finally, the model undergoes fine-tuning guided by a joint-level distorter to ensure optimal performance. To validate the effectiveness of ScreenMark, we compiled a dataset comprising 100,000 screenshots from various devices and resolutions. Extensive experiments across different datasets confirm the method’s superior robustness, imperceptibility, and practical applicability.
@@ Introduction
With the continuous advancement of the Internet and computer technologies, an increasing amount of information is presented in the form of Visual Screen Content (VSC), including images, videos, texts, web pages, windows, and more. In the context of users’ personal computers, VSC is displayed on screens without exception. From a visual perspective, this mode of presentation is the most readily acceptable form of expression. At the same time, this also means that the VSC will be easily leaked.
@@
To address the limitations in protection scope and response times encountered by current watermarking techniques for screen content, this paper introduces a three-stage watermarking framework specifically designed for screen content. Inspired by the diffusion model(Ho, Jain, and Abbeel 2020; Dhariwal and Nichol 2021), which diffuses a regular image with noise to generate an irregular image, this framework adopts a novel approach. The diffusion process can be reversed through denoising, restoring the original image. This suggests directly integrating irregular watermark patterns obtained from the diffusion of regular watermark information with the screen content, rather than embedding information into the protected multimedia carrier using an encoder, as is common in traditional watermarking frameworks. This is particularly important for screen content protection because it allows real-time and cost-effective integration of watermark patterns with screen content. It offers a protection method that is unrestricted by scope and response time, integrating securely and naturally with screen content.
@@
Initially, we generate a batch of regular watermark messages ${I_{w}}$ , with batch size ${N_{0}}$ and information length ${L}$. Subsequently, ${I_{w}}$ is subjected to diffusion processing within the ${M_{D}}$, resulting in an irregular watermark pattern ${P_{w}}$.Upon acquiring ${P_{w}}$, each pattern in batch undergoes parallel image-level distortion processing by ${\mathcal{D}_{I}^{k}\in D_{I}}$, where ${k\in\{1,2,\ldots,N_{0}\}}$, yielding the distorted watermark pattern ${P_{d}}$. Finally, the reverse-processed watermark information ${I_{r}}$ is retrieved through the ${M_{R}}$.
@@
Table 1: Bit accuracy rate(BAR,%) on different image-level attacks
@@
Table 2: Bit accuracy rate(BAR,%) on different pixel-level attacks
@@
Figure 4: The Loss vs. Epoch of different training strategy
@@
Digital watermarking for VSC must be capable of real-time protection, as screenshot commands can be scripted to achieve millisecond-level interception. Existing single-modal watermarks can not embed in dynamic screen content in real time. We validated the responsiveness of SOTAs to VSC changes. As shown in Tab.4, our watermark, fused directly to the screen, achieves a reaction time of 0 milliseconds, outperforming SOTAs. Notably, the time here is not the execution time of program, but the reaction time to reload the watermark information to the new VSC when it changes.
Table 4: Average reaction time of SOTAs in VSC scenarios
@@ Conclusion
In this paper, we propose ScreenMark, a robust deep-learning-based robust watermarking scheme for arbitrary VSC protection, using a three-stage progressive watermarking training strategy. The message diffuser and message reverser facilitate the transformation between regular watermarking information and irregular watermarking patterns. The alpha-fusion rendering module integrates these patterns into VSCs of any resolution, while the screen decoder extracts the watermark information from distorted watermarked screenshots. We built a dataset with 100,000 screenshots from various devices and resolutions. Extensive experiments demonstrate ScreenMark’s effectiveness in robustness, imperceptibility, and practical applicability.

### v3

@@ Abstract
Digital watermarking has shown its effectiveness in protecting multimedia content. However, existing watermarking is predominantly tailored for specific media types, rendering them less effective for the protection of content displayed on computer screens, which is often multi-modal and dynamic. Visual Screen Content (VSC), is particularly susceptible to theft and leakage through screenshots, a vulnerability that current watermarking methods fail to adequately address. To address these challenges, we propose ScreenMark, a robust and practical watermarking method designed specifically for arbitrary VSC protection. ScreenMark utilizes a three-stage progressive watermarking framework. Initially, inspired by diffusion principles, we initialize the mutual transformation between regular watermark information and irregular watermark patterns. Subsequently, these patterns are integrated with screen content using a pre-multiplication alpha blending technique, supported by a pre-trained screen decoder for accurate watermark retrieval. The progressively complex distorter enhances the robustness of the watermark in real-world screenshot scenarios. Finally, the model undergoes fine-tuning guided by a joint-level distorter to ensure optimal performance. To validate the effectiveness of ScreenMark, we compiled a dataset comprising 100,000 screenshots from various devices and resolutions. Extensive experiments on different datasets confirm the superior robustness, imperceptibility, and practical applicability of the method.
@@ Introduction
With the continuous advancement of the Internet and computer technologies, an increasing amount of information is presented in the form of Visual Screen Content (VSC), including images, videos, texts, web pages, windows, and more. In the context of users’ personal computers, VSC is displayed on screens without exception. From a visual perspective, this mode of presentation is the most readily acceptable form of expression. At the same time, this also means that the VSC can easily be leaked.
@@
To address the limitations in protection scope and response times encountered by current watermarking techniques for screen content, this paper introduces a three-stage watermarking framework specifically designed for screen content. Inspired by the diffusion model(Dhariwal and Nichol 2021), which diffuses a regular image with noise to generate an irregular image, this framework adopts a novel approach. The diffusion process can be reversed through denoising, restoring the original image. This suggests directly integrating irregular watermark patterns obtained from the diffusion of regular watermark information with the screen content, rather than embedding information into the protected multimedia carrier using an encoder, as is common in traditional watermarking frameworks. This is particularly important for screen content protection because it allows real-time and cost-effective integration of watermark patterns with screen content. It offers a protection method that is unrestricted by scope and response time, integrating securely and naturally with screen content.
@@
Initially, we generate a batch of regular watermark messages ${I_{w}}$ , with batch size ${N_{0}}$ and length of information ${L}$. Subsequently, ${I_{w}}$ is subjected to diffusion processing within ${M_{D}}$, resulting in an irregular watermark pattern ${P_{w}}$. Upon acquisition of ${P_{w}}$, each pattern in batch undergoes parallel image-level distortion processing by ${\mathcal{D}_{I}^{k}\in D_{I}}$, where ${k\in\{1,2,\ldots,N_{0}\}}$, yielding the distorted watermark pattern ${P_{d}}$. Finally, the reverse-processed watermark information ${I_{r}}$ is recovered through ${M_{R}}$.
@@
Table 1: Bit accuracy rate(BAR,%) on different image-level attacks
@@
Table 2: Bit accuracy rate(BAR,%) on different pixel-level attacks
@@
Figure 4: The Loss & Acc vs. Epoch of training strategy
@@
Digital watermarking for VSC must be capable of real-time protection, as screen capture commands can be scripted to achieve millisecond level interception. Existing single-modal watermarks can not embed in dynamic screen content in real time. We validated the responsiveness of SOTAs to VSC changes. As shown in Tab.4, our watermark, fused directly to the screen, achieves a reaction time of 0 milliseconds, outperforming SOTAs. In particular, here the time is not the execution time of program, but the reaction time to reload the watermark information to the new VSC when it changes.
@@
Table 4: Average reaction time of SOTAs in VSC scenarios
@@ Conclusion
In this paper, we propose ScreenMark, a robust deep learning-based robust watermarking scheme for arbitrary VSC protection, using a three-stage progressive watermarking training strategy. The message diffuser and message reverser facilitate the transformation between regular watermarking information and irregular watermarking patterns. The alpha-fusion rendering module integrates these patterns into VSCs of any resolution, while the screen decoder extracts the watermark information from distorted watermarked screenshots. We built a dataset with 100,000 screenshots from various devices and resolutions. Extensive experiments demonstrate the effectiveness of ScreenMark in robustness, imperceptibility, and practical applicability.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants U20B2051, U22B2047, 62450067, 62072114.